OpenStudy (anonymous):

Solve x – 9 ≥ -7(x + 2). Write your solution using set builder notation. Please show all of your work.

OpenStudy (anonymous):

\[x-9 \ge -7x-14 \implies 8x \ge -5; x \ge -\frac{5}{8} \implies S= \left\{ x| x \in [-5/8, \infty) \right\}\] Where S is the solution set.

OpenStudy (anonymous):



OpenStudy (anonymous):

The number on the top is 5, sorry about that.